The molar volume of a cetain solid is 159.5 cm3 mol-1 at 1.00 atm and 350.05 K, its melting temperature. The molar volume of the liquid at this temperature and pressure is 161.8 cm3 mol-1. At 100 atm the melting temperature changes to 350.56 K. Calculate the entropy of fusion of the solid .
Answer #1

we know that the entropy of fusion can be calculated by calsuius equation ,

dp / dT = \Delta​Sfus / \DeltaV

putting all the values , dp = 100 atm - 1 atm = 99 atm , dT = 350.56 - 350.05 = 0.51 K

and \Delta​V = 161.8 - 159.5 = 2.3 cm3mol-1

therefore

99 atm / 0.51 K = \Delta​Sfus / 2.3 cm3mol-1

\Delta​Sfus​ = 446.47 L-atm/K-mol

\Delta​Sfus​ = 45212.57 J/K-mol